# A note on thermodynamics

### v1

@@ Abstract
The process of creating goods and services, measured by their value, is considered as a process of creating complexity. This makes it possible to consider the production system as an open thermodynamic system, and to develop a simple heuristic model for the production process. The model includes three production factors: the index of complexity of production equipment (physical capital $K$), human activity (labour $L$), and the substitutive capacity of equipment (substitutive work $P$). The latter is a contribution to economic theory from the thermodynamic approach, which also requires the introduction of technological characteristics of production equipment, such as labor requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labor and energy required to operate production equipment. By applying thermodynamic principles to the theory of production, we can understand how labour can be replaced by capital, and derive the production function in four equivalent but different formulations. Two of them are known and used by economists for interpretation the production phenomena; the thermodynamic approach gives some foundation for economic theory. The production function allows an unambiguously decompose of the growth rate of output according to the growth rates of production factors and technological level. The introduction of substitute work as a factor of production and technological features of capital expands planning and analyse of production processes.
@@
capital, labour, substitutive work, technological coefficients, production function
@@ Introduction
Any production of consumables (food, clothes, buildings, transport, sewage systems, household appliances, machinery and so on) contains processes that convert natural resources into products for human usage. You can observe clay being turned into pots; or clay, sand and stone being transformed into buildings. Raw materials such as ore are turned into airplanes, which fly in the sky. This approach allows us to view the products we create as a result of transforming ’wild’ material into ’usable’ forms through work done by humans or machines (using energy from external sources such as wind, water, or coal). Any production system is being realized as a set of appliances for the matter transformations, and the system can be considered to be an open thermodynamic system. The simplest schematization of the production procces is shown in Fig. 1.
The aim of the production process is to obtain useful output $Y$, which is a monetory estimate of market value of created things. The production processes can be considered as processes of creation of value, and researchers have been engaged in finding the origins of this property, so called production factors. Historically, among the three factors listed in the input on Fig. 1, the aggregate human efforts — labour $L$ — was chosen by classical political economists [1, 2] as the sole source of value. That was a perfect solution until machines began to dominate workers’ efforts in production processes. To solve the problems that arise, the researchers incorporate another factor into their description. It was the amount of production equipment, also known as physical capital $K$, measured by its value, that was chosen as an aditional production factor.
Conventionally, the output of a production process, $Y$ is now understood as the result of activity of these two factors [3]. The first factor is labour efforts, $L$, which can be quantified, for example, in working hours per year. The second factor is physical capital, $K$, which represents the quantity of equipment used in production, so that the production of value
@@
The relation describes the effect of substitution; this means that the two production factors could substitute for each other. In other words, productive power is prescribed not only to labour, but also to the installed equipment; both of them appear to be the sources of value. The researchers are using the various specifications of function (1), adding some correcting assumptions, if necessary [3] .
Looking at Fig. 1, we are surprised; why was the amount of production equipment $K$, but not its work $P$, chosen as an argument for the production function. Indeed, how can labour be substituted by the amount of production equipment? One has to at least understand the mechanism of this substitution. Moreover, relation (1) could lead to some contradictions that are described in the literature [4].
In this paper, we propose considering the approach based on the concepts and principles of thermodynamics. The output of the production processes $Y$ is conventionally interpreted as market value estimated by consumers; however, this does not mean that it is a subjective quantity. The objective content of output appears to be complexity; the production of consumable goods can be interpreted as a process of creating useful complexity within the human environment; production activity moves boundaries between the artificial and natural environment. The created complexity $Y$ can be interpreted also as entropy that is diminishing in the human environment [5], which allows us to state that the natural measure of value is energy [5, 6]. Creation of complexity requires work that could be done by labor efforts, $L$, and/or machine work, $P$. Therefore, according to the thermodynamic laws, we can write
@@
Functions (1) and (2) represent two different approaches to describing production processes. There is a contradiction between these approaches, and a clear desire exists to find a unified description, which is the main goal of this paper. We believe that a thermodynamic approach can be useful in formulating the theory. Based on a simple diagram shown in Fig.1, we will develop in the next section a simple model that will allow us to determine the production function of a process in three forms. One form corresponds to relation (1), which describes the phenomenon of substitution, a well-known phenomena.
@@
To determine specific forms of functions (1) and (2), we turn to the simple model of the production process, shown in Figure 1. The assembly of production equipment, that is a fixed embodiment of technology, commonly [3] characterised by physical capital $K$, measured in units of value [3] , can be considered as an open thermodynamic system. From a thermodynamic perspective, the variable $K$ is a characteristic of a non-equilibrium metha-stable state of the system, which appears to be the index of complexity. The production equipment is an essential part of modern enterprises, but it remains inactive until efforts by workers and energy input from external sources have been applied. The installed equipment determines how much labour $L$ and substitutive work $P$ will be required for the production process before it is activated. We assume that these factors of production are measured in energy units.
@@
If the quantity is less than one, it means that labor-saving or energy-saving technologies have been used at that moment of time. In a stationary case, the quantities do not change with time.
@@
From a thermodynamic point of view, the true and only sources of value (as relatives of entropy) are labor expenses $L$ and the work of external energy sources $P$. Human efforts are applied directly, whereas substitutive work arises from the prior exploration of natural laws, the development of technological systems, and the manufacturing of production machinery, among other things. The interpretation of the production process allows us to view the output $Y$ as a function of two variables: the activity of workers — labour $L$ — and the work done by production equipment - substitutive work $P$, which depends, according to equation (5), on the amount of installed equipment $K$, so that we can write the following function
@@
As the amounts of work, the production factors are interchangeable and, in this sense, they are equivalent in all respect. The Smith-Marx theory of labor value is supplemented by the law of substitution, which says that work of third-party forces of nature via production equipment replaces human efforts. Labor functions as labor plus work of equipment in complex in the production of value.
@@
The index $\alpha$ — the technological index — in equation (8) determines the degree of substitution of labour with productive energy. It was shown in [8, Chapter 6; 9], considering the dynamics of the production factors, that the index $\alpha$ in equation (8) has the same form as it is defined by the formula (6); the level of technology development is included in the law of value production. Note that the labor $L$ and substitutive work of external energy sources $P$ is measured in energy units. Relation (8) was confirmed in many details in the author’s previous works [8, 9].
Relation (8) describes the effect of substitution labor efforts with machine work. Economists usually prefer to talk about substituting labor with capital [3]. Relations (5) allow us to represent the production of value (8) in terms of the variables $L$ and $K$. When we choose these variables, the resulting production function is as follows
@@
The recorded relation is a specific form of function (1), and represents the law of labor substitution by capital. An increase in output can be achieved through increasing relative labor expenses, or the relative amount of capital. The substitution coefficient depends on technological characteristics of production equipment. An analysis of substitution mechanisms shows that workers’ efforts to produce goods are substituted by the work of the production equipment, not by equipment itself.
@@
One of the relations (5) shows the amount of human effort for the presence of production equipment (physical capital $K$), so that the relations (9) and (10) are reduced to the form
@@
This relation gives us the reason to say that labor remains, to use the words of Adam Smith, ”the only universal, as well as the only accurate measure of value, or the only standard by which we can compare the values of various goods at all times and in all places”. The law of the value production appears to be a nonlinear, generally speaking, function.
@@
Three forms of production function give rise to three different interpretations of processes of value production . Economists usually prefer to interpret the situations in terms of labor and capital [3]. Within the framework of this interpretation, the concept of the production power of capital has been formed, and the idea of capital productivity has become ingrained in the public consciousness. This applies not only to physical capital, but also to financial capital in various forms. For example, if we own shares in a company, we receive dividends; if we deposit money in a bank, we earn interest. Shares and money represent capital in its broadest sense.
@@
In the paper, an interpretation of production process as process of creating complexity was proposed. In order to explain and describe increased labor productivity through mechanization, we need to take into account three fundamental factors of production: capital ($K$), labor ($L$), and substitute work ($P$). The latter is a contribution to economic theory from the thermodynamic approach, which also requires the introduction of technological characteristics of production equipment, such as labor requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labor and energy required to operate production equipment. By invoking thermodynamic principles in the theory of production processes, we can explain how labour is replaced by capital in production and derive a production function with four equivalent formulations (8) - (11). One can be satisfied that two forms (9) and (10) were known and were used by the economists for interpretation the production phenomena. In other words, the thermodynamic approach gives some foundation for relations of economic theory. The introduction of technological features of capital and substitute work as a factor of production expands the ability to plan and analyze the production. According to the theory, when describing the production of goods, qualitative technological attributes of capital matter more than its overall amount.
In conclusion, we should emphasize the fundamental importance of the law of value production. This law plays a crucial role in shaping the interpretation of production processes, and it serves as a foundation for determining how social goods are distributed. In essence, this law holds a fundamental position within the economic system and therefore deserves recognition as an essential economic principle. Finally, this question concerns the correct understanding of observable social phenomena.

### v2

# Thermodynamics

@@ Abstract
The process of creating goods and services, measured by their value, is considered a process of creating complexity. This allows us to consider the production system as an open thermodynamic system, and to develop a simple heuristic model of the production process. The model includes three production factors: the index of complexity of production equipment (physical capital $K$), human activity (labour $L$), and the substitutive capacity of equipment (substitutive work $P$). The latter is a contribution to the theory of production from the thermodynamic approach, which also involves introducing technological characteristics of production equipment, such as labour requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labour and energy required to operate production equipment. By applying thermodynamic principles, we can understand how labour can be replaced by capital and derive the production function with four different formulations. Two of them are known and used by researchers for interpretation the production phenomena; the thermodynamic approach provides some foundation for economic theory, allowing us to decompose unambiguously the growth rate of output over technological level and the growth rates of production factors. Introducing substitute work as a factor of production and technological characteristics of capital expands our ability to plan and analyse production processes.
@@
capital, , substitutive work, technological coefficients, production function
@@ Introduction
Any production of consumables (food, clothes, buildings, transport, sewage systems, household appliances, machinery and so on) contains processes that convert natural resources into products for human usage. You can observe clay being turned into pots; or clay, sand and stone being transformed into buildings. Raw materials such as ore are turned into air plane, which fly in the sky. These pictures allow us to view the products that are created as a result of transforming ’wild’ materials into ’usable’ forms with the work done by humans or machines (using energy from external sources such as wind, water, or coal). Any production system is being realized as a set of appliances for the matter transformations, and the system itself can be considered to be an open thermodynamic system. The simplest schematization of the production process is shown in Fig. 1.
The aim of the production process is to obtain useful output $Y$, which is a monetary estimate of market value of created things. The production processes can be considered as processes of creation of value, and many researchers have been engaged in finding the origins of the wealth, so called production factors. Historically, among the three factors listed in the input on Fig. 1, the aggregate human efforts — labour $L$ — was chosen by classical political economists [1, 2] as the sole source of value. That was a perfect solution until machines began to dominate workers’ efforts in production processes. To solve the problems that arise in the understanding, the researchers incorporate another factor into their description. It was the amount of production equipment, measured by its value, also known as physical capital $K$, that was chosen as an additional production factor. Conventionally, the output of a production process, $Y$ is now understood as the result of activity of these two factors [3], so that the production of value can bt expresed as
@@
The first factor is labour efforts, $L$, which can be quantified, for example, in working hours per year. The second factor is physical capital, $K$, which represents the quantity of equipment used in production.
The relation (1) describes the effect of substitution; this means that the two production factors could substitute for each other. In other words, the property to create value is prescribed not only to labour, but also to the installed equipment; both of them appear to be the sources of value. The researchers are using function (1) in various specifications, adding some correcting assumptions, if necessary [3] . However, one can be puzzled: why was the amount of production equipment $K$, but not its work $P$, chosen as an argument for the production function, which seems to be natural, looking at Fig. 1. Indeed, how can labour be substituted by the amount of production equipment? One has to at least understand the mechanism of this substitution. Moreover, relation (1) could lead to some contradictions that are described in the literature [4].
In this paper, we propose considering the approach based on the concepts and principles of thermodynamics, interpreting the production system to be an open system, the non-equilibrium state of which is characterized with internal variable — physical capital $K$ . The output of the production processes $Y$ is conventionally interpreted as market value estimated by consumers; however, this does not mean that it is a subjective quantity. The objective content of output appears to be complexity; the production of consumable goods can be interpreted as a process of creating useful complexity within the human environment; production activity moves boundaries between the artificial and natural environment. The created complexity $Y$ can be interpreted also as entropy that is diminishing in the human environment [5], which allows us to state that the natural measure of value is energy [5, 6]. Creation of complexity requires work that could be done by labour efforts, $L$, and/or machine work, $P$. Therefore, according to the thermodynamic laws, we can write in general form
@@
Functions (1) and (2) represent two different approaches to describing production processes. There is a clear desire to find a unified description, which is the main goal of this paper. We believe that a thermodynamic approach can be useful in formulating the theory. Based on a simple diagram shown in Fig.1, we will develop in the next section a simple model that will allow us to determine the production function of a process in four forms. One form corresponds to relation (1), which describes the phenomenon of substitution, a well-known phenomena.
@@
To determine specific forms of functions (1) and (2), we turn to the simple model of the production process, shown in Figure 1. The assembly of production equipment, that is a fixed embodiment of technology, commonly [3] characterised by physical capital $K$, measured in units of value [3] , can be considered as an open thermodynamic system. From a thermodynamic perspective, the variable $K$ is a characteristic of a non-equilibrium metha-stable state of the system, which can be considered as the index of complexity. The production equipment is an essential part of modern enterprises, but it remains inactive until efforts by workers and energy input from external sources have been applied. The installed equipment determines how much labour $L$ and substitutive work $P$ will be required for the production process before it is activated. We assume that these factors of production are measured in energy units.
@@
If the quantity is less than one, it means that labour-saving or energy-saving technologies have been used at that moment of time. In a stationary case, the quantities do not change with time.
@@
From a thermodynamic point of view, the true and only sources of value (as relatives of entropy) are labour expenses $L$ and the work of external energy sources $P$. Human efforts are applied directly, whereas substitutive work arises from the prior exploration of natural laws, the development of technological systems, and the manufacturing of production machinery, among other things. The interpretation of the production process allows us to view the output $Y$ as a function of two variables: the activity of workers — labour $L$ — and the work done by production equipment - substitutive work $P$, which depends, according to equation (5), on the amount of installed equipment $K$, so that we can write the following function
@@
As the amounts of work, the production factors are interchangeable and, in this sense, they are equivalent in all respect. The Smith-Marx theory of labour value is supplemented by the law of substitution, which says that work of third-party forces of nature via production equipment replaces human efforts. The labour acts as labour plus work of equipment in complex in the production of value.
@@
The index $\alpha$ — the technological index — in equation (8) determines the degree of substitution of labour with productive energy. It was shown in [8, Chapter 6; 9], considering the dynamics of the production factors, that the index $\alpha$ in equation (8) has the same form as it is defined by the formula (6); the level of technology development is included in the law of value production. Note that the labour $L$ and substitutive work of external energy sources $P$ is measured in energy units. Relation (8) was confirmed in many details in the author’s previous works [8, 9].
Relation (8) describes the effect of substitution labour efforts with machine work. Economists usually prefer to talk about substituting labour with capital [3]. Relations (5) allow us to represent the production of value (8) in terms of the variables $L$ and $K$. When we choose these variables, the resulting production function is as follows
@@
The recorded relation is a specific form of function (1), and represents the law of labour substitution by capital. An increase in output can be achieved through increasing relative labour expenses, or the relative amount of capital. The substitution coefficient depends on technological characteristics of production equipment. An analysis of substitution mechanisms shows that workers’ efforts to produce goods are substituted by the work of the production equipment, not by equipment itself.
@@
One of the relations (5) shows the amount of human effort for the presence of production equipment (physical capital $K$), so that the relations (9) and (10) can be are represented in the form
@@
This relation gives us the reason to say that labour remains, to use the words of Adam Smith, ”the only universal, as well as the only accurate measure of value, or the only standard by which we can compare the values of various goods at all times and in all places”. The law of the value production appears to be a non-linear, generally speaking, function.
@@
Three forms of production function give rise to three different interpretations of processes of value production . Economists usually prefer to interpret the situations in terms of labour and capital [3]. Within the framework of this interpretation, the concept of the production power of capital has been formed, and the idea of capital productivity has become ingrained in the public consciousness. This applies not only to physical capital, but also to financial capital in various forms. For example, if we own shares in a company, we receive dividends; if we deposit money in a bank, we earn interest. Shares and money represent capital in its broadest sense.
@@
Decomposition of the output growth rate
Expressions (8) — (11) demonstrate that, according to empirical data, at any moment the value of manufactured products $Y$, measured in monetary units of constant purchasing power, is determined by the technological characteristics of production equipment: energy requirement $\overline{\varepsilon}$ and labour requirement $\overline{\lambda}$, also as by the aggregated amounts of social resources (production factors) used: physical capital $K$, direct efforts of people $L$ (physical and mental, we cannot separate them) and the substitute work of machines $P$ that is obtained with the help of natural energy sources such as wind, water, coal and so on. The production factors are determined by the availability of social resources and technological capabilities (the technological characteristics) of the production equipment and are, generally speaking, functions of time.
The influence of production factors and technological characteristics on economic growth is clearly demonstrated when calculating the growth rate of output $Y$, for which it is necessary to write down the full differential of output, referring to the relations (8) – (11). For example, referring to the expression (8) and (11) allows one to write down the relations
$\frac{1}{Y}\frac{\rm dY}{\rm dt}=(1-\alpha)\frac{1}{L}\frac{\rm dL}{\rm dt}+\alpha\frac{1}{P}\frac{\rm dP}{\rm dt}+\ln\left(\frac{L_{0}}{L}\frac{P}{P_{0}}\right)\frac{\rm d\alpha}{\rm dt}.$ (14)
$\frac{1}{Y}\frac{\rm dY}{\rm dt}=\frac{1}{\overline{\lambda}}\left[\frac{1}{L}\frac{\rm dL}{\rm dt}-\ln\left(\frac{L}{L_{0}}\right)\frac{1}{\overline{\lambda}}\frac{\rm d\overline{\lambda}}{\rm dt}\right].$ (15)
The growth rate of output is determined by the growth rate of production factors, while the ’structure’ of the production system remain unchanged, and by the technological characteristics of the equipment (the latter terms are on the right side of equations 14 and 15). With a constant value of the technological index, an increase in output is associated with an increase in the use of production factors with their constant characteristics, $P/L=const$, which is defined as extensive economic growth. The last term in formulae (14) and (15) is directly related to structural and/or technological changes in the production system, that is, to the evolution of the production system itself. As the technological indexes increase, the $P/L$ ratio also increases, and the case of intensive economic growth is realized.
To exclude the presence of zero values in the right-hand sides of ratios (14) and (15), which indicates that a change in technological characteristics with constant resource consumption ($L=L_{o},P=P_{o}$) does not lead to any change in output, the last terms in these ratios can be written in a different form when using an expression that is valid in linear approximation by increment
$\ln\left(\frac{L}{L_{0}}\right)\approx\frac{1}{L}\Delta L.$
The relations (14) and (15) now take the form
$\frac{1}{Y}\frac{\rm dY}{\rm dt}=(1-\alpha-\Delta\alpha)\frac{1}{L}\frac{\rm dL}{\rm dt}+(\alpha+\Delta\alpha)\frac{1}{P}\frac{\rm dP}{\rm dt}$ (16)
$\frac{1}{Y}\frac{\rm dY}{\rm dt}=\frac{1}{\overline{\lambda}}\left[1-\frac{1}{\overline{\lambda}}\,\Delta{\overline{\lambda}}\right]\frac{1}{L}\frac{\rm dL}{\rm dt}.$ (17)
The latter formula is convenient for analysing and evaluating the two policy options for increasing output: direct labour expenses can be increased (additional labour is attracted) or technology and third-party energy consumption can be improved. The technological coefficient (labour requirement) $\overline{\lambda}$ decreases with increasing machine work, and with a decrease equal to the coefficient value, the effect of changing the technological characteristics of the system is compared with the effect of a direct increase in labour expenses.
For the actual decomposition of the output growth rate, time series on output $Y$ and production factors are required. The decomposition of the output growth rate is unambiguous and does not require additional information: there are no arbitrary parameters included in the theory.
@@
The thermodynamic approach helps us to solve one of the fundamental economic problem — the problem of origin of the wealth in the epoch, when different machine and appliances are widely used. Up to now the existing formulae for the calculation of output do not contain the appropriate technological characteristics of production equipment. The output of a production system is determined through production factors in various ways; in one of the simplest and most commonly used case, for example , a production function can be written [3] in the terms of capital ($K$) and labour ($L$) as
$Y=Y_{0}A(t)\,\frac{L}{L_{0}}\left(\frac{L_{0}}{L}\frac{K}{K_{0}}\right)^{a},\quad 0<a<1.$ (18)
This relation represents the Cobb-Douglas function with an additional empirical multiplier $A(t)$, which is introduced, as evidence that there is something what we do not know. Under proper choice of the time-dependent multiplier $A(t)$, formula (18) allows one to achieve a complete description of the time dependence of output. The sense of the adjustment multiplier $A(t)$, which is assumed to describe exogenous technological progress, remains unexplained, despite various interpretations and modifications of the production function. All these raises doubts about the adequacy of such a description. The comparison with similar presentation of the production function (9) allows us to calculate the exogenous technological progress in function (18) , but the matter is that the form of the empirical function (18) is inappropriate. The complex construction (18) could arise only with an initial misunderstanding of the essence of the phenomena of substitution labour with capital.
In this paper, we have calculated the law of the value production, considering the assembly of production equipment as an open thermodynamic system and the production processes as processes of creating complexity. In order to explain and describe increased labour productivity through mechanization, we need to take into account three fundamental factors of production: physical capital ($K$), labour ($L$), and substitute work ($P$). The latter is a contribution to the theory of production from the thermodynamic approach, which also requires the introduction of technological characteristics of production equipment, such as labour requirement ($\overline{\lambda}$) and energy requirement ($\overline{\varepsilon}$), which indicate the amounts of labour and energy required to operate production equipment. By invoking thermodynamic principles in the theory of production processes, we can explain how labour is replaced by capital in production and derive a production function with four equivalent formulations (8) - (11). One can be satisfied that two forms (9) and (10) were known and were used by researches for interpretation the production phenomena. In other words, the thermodynamic approach gives the solid foundation for relations of economic production theory. According to the theory, when describing the production of goods, qualitative technological attributes of capital matter more than its overall amount. The introduction of substitute work as a factor of production and the technological features of capital expands the ability to plan and analyse the economic processes.
In conclusion, we should emphasize the fundamental importance and fundamental nature of derived results. These is also evidenced by the fact that earlier every small advance in understanding the relationship between output and production factors has been rewarded with the so-called Nobel Prize in Economics. The law of value production plays a crucial role in shaping the interpretation of production processes, and it serves as a foundation for determining how social goods are distributed. In essence, this law holds a fundamental position within the social system and therefore deserves recognition as an essential economic principle. Finally, this question concerns the correct understanding of observable social phenomena.